A ray of light passes from water through a plate of crown glass and on to air. The angle of incidence in the water is 40.0 °.

What then is the angle of refraction in the glass?
1
Expert's answer
2020-06-01T14:19:27-0400

We can use Snell's law to calculate the angle of refraction "\theta_2" in glass. We have two media: water with n1=1.33 and glass with n2=1.655. Then


"\frac{\text{sin}\theta_1}{\text{sin}\theta_2}=\frac{n_2}{n_1},\\
\theta_2=\text{arcsin}\bigg(\text{sin}\theta_1\frac{n_1}{n_2}\bigg)=\\\space\\

\theta_2=\text{arcsin}\bigg(\text{sin}40°\frac{1.33}{1.655}\bigg)=31°."